A supplemental angle is 2 times the compliment plus 40, what is the angle

Answer by macston(5194)

It is unclear which situation you mean.
.
x=the angle
.
For (two times the complement) plus 40.
180-x=2(90-x)+40 or
180-x=180-2x+40
-x=-2x+40
x=40
.
-OR-
.
For two times (the complement plus 40).
180-x=2(90-x+40)
180-x=180-2x+80
-x=-2x+80
x=80